# Statistical Inference Leveraging Synthetic Data with Distribution-Free Guarantees

### v1

@@ C.2 Conformal risk control for protein structure prediction
Figure˜13 visualizes protein T1078, for which AlphaFold produces a highly accurate predicted structure. Figures˜11(a) and 11(b) present the resulting prediction sets obtained by OnlyReal and GESPI, respectively. Panel (1) presents the predicted structure, with residues abstained on marked in red and accepted residues in blue. Both methods obtain risk equal to 0; however, OnlyReal conservatively abstains from most residues ($\sim 91\%$), while GESPI abstains from significantly fewer ($\sim 48\%$), demonstrating the benefit of leveraging synthetic data. Panel (2) shows the predicted structure (red and blue) aligned with the true structure (gray), illustrating that most of the protein is accurately predicted. Panel (3) highlights a small subset of residues whose predicted structure is inaccurate; both methods abstain. Lastly, panel (4) shows a well-predicted region where OnlyReal unnecessarily abstains, while GESPI correctly accepts the residues.
@@
Figure 11: Visualization of protein structure prediction with error rate control. Panels show protein T1078 predictions with residues abstained on by (a) OnlyReal and (b) GESPI methods. Red: residues abstained on; Blue: accepted residues. Gray: real experimental structure, aligned with AlphaFold2 predicted structure. Quantitative results {abstention ratio, risk}: OnlyReal – $\approx 91\%$, $0\%$}; GESPI – $\approx 48\%$, $0\%$}.
Figure˜12 complements Figure˜3 in the main manuscript by showing the chosen pLDDT thresholds, $\hat{\lambda}$, for $\alpha=5\%$ and $10\%$. The figure demonstrates that the selected thresholds vary with $\alpha$ and differ across methods: OnlyReal conservatively chooses higher thresholds, while our GESPI procedure selects lower thresholds, which, as shown in Figure˜3, result in risk closer to the nominal $\alpha$ level.
@@
Figure 12: Chosen pLDDT thresholds, $\hat{\lambda}$, for $\alpha=5\%$ and $10\%$, complementing the results shown in Figure˜3.
@@
Figure˜13 presents the performance of the conformal risk control methods at target level $\alpha=15\%$. The results show a similar trend to that in Figure˜3. In particular, OnlyReal conservatively controls the risk and leads to a higher and more variable abstention rate. OnlySynth, which relies on approximate prediction errors, does not provide valid risk control guarantees. In contrast, our proposed method, GESPI, achieves risk close to the nominal $\alpha$ level with a lower abstention rate.
@@
Figure 13: Performance comparisons for protein structure prediction with error rate control. Conformal risk control methods applied at target level $\alpha=15\%$. Left: average risk (fraction of residues with error >3>3Å). Middle: average abstention rate. Right: selected pLDDT threshold $\hat{\lambda}$.
C.3 Single and multiple hypothesis testing for outlier detection
We extend the experiments from Section˜4.3 by rerunning them with a different trimming proportion. Given that the true fraction of outliers in the contaminated reference set is generally unknown, we evaluate the performance under a smaller trimming rate, reducing it from $q=5\%$ (main manuscript) to $q=2.5\%$, i.e., removing fewer datapoints from the contaminated reference set.
As shown in Figure˜14, the results follow a similar trend as in the main manuscript. OnlySynth produces error rates both above and below the nominal level, as it does not provide any error rate control guarantees. OnlyReal conservatively controls the error metric at the target level, and GESPI achieves error rates close to the target while improving power, approaching the performance of the Oracle.
@@ C.3 Single and multiple hypothesis testing for outlier detection
Figure 14: Performance comparisons for outlier detection. Evaluated on three datasets: (1) shuttle, (2) credit-card, (3) KDDCup99. Left two panels: single-outlier case ($\alpha=2\%$, $\varepsilon=1\%$). Right two panels: multiple-outlier case ($\alpha=15\%$, $\varepsilon=10\%$). Same setup as in Figure˜5, but with $q=2.5\%$ trimming proportion.
C.4 Hypothesis testing for mechanistic interpretability of a Vision Transformer model
@@
In this section, we consider a task related to the mechanistic interpretability of vision transformers. We aim to test whether a specific attention head exhibits a functional role—for example, detecting shapes, animals, or spatial locations. Concretely, we test whether the activation of a given head is stronger when the object it is hypothesized to detect is present in the image compared to when it is absent, which would provide evidence that the head indeed serves this role. As a case study, we examine attention head L22H6 in CLIP ViT-L/14 (Radford et al., 2021), which has been previously reported to strongly respond to images with animals (Gandelsman et al., 2024).
@@
We use ImageNet images as the real dataset, partitioned into animals and non-animals classes. For the synthetic dataset, we generate images using FLUX.1 (Labs, 2024), also separated into animals and non-animals. Details for each dataset are as follows:
@@
  • FLUX.1 (synthetic) (Labs, 2024): For each class, we generate 2,0002{,}000 images using the FluxPipeline from diffusers with mixed precision (float16) and $50$ inference steps. Prompts follow the CLIP-style template “A photo of a {class name}” (Radford et al., 2021), where {class name} is the corresponding ImageNet label. The resulting synthetic images are labeled animals / non-animals based on the originating class.
@@
Intuitively, $H^{l,h}_{i}(X)$ measures how strongly token $i$ aligns with the animal text direction. See an example visualized as a spatial heatmap in Figure 15.
Figure 15: Left: input image $X$ of a dog. Right: spatial heatmap $H^{22,6}(X)$ from layer 6, head 22, obtained by projecting token contributions onto $M_{\text{text}}(\text{``a photo of an animal''})$.
@@
Formulation of the null and alternative. We compute the activation score $s^{l,h}(X)$ (8) on two disjoint image sets, $\mathcal{D}_{\text{animals}}$ and $\mathcal{D}_{\text{non-animals}}$, and test whether head $(l,h)$ is more active on animals. Let $P^{(l,h)}_{\text{animals}}$ and $P^{(l,h)}_{\text{non-animals}}$ denote the population distributions of activation scores of head $(l,h)$ over animals and non-animals, respectively. Let $\mu^{(l,h)}_{\text{animals}}$ and $\mu^{(l,h)}_{\text{non-animals}}$ denote the expected values/means of these populations. We test the null hypothesis that these two distributions are equal against the one-sided alternative that the mean activation score over the animal distribution is larger:
$\mathcal{H}_{0}:P^{(l,h)}_{\text{animals}}=P^{(l,h)}_{\text{non-animals}}\text{ versus }\mathcal{H}_{1}:\mu^{(l,h)}_{\text{animals}}-\mu^{(l,h)}_{\text{non-animals}}>0.$
@@
To set the critical value, we use a standard permutation approach, which is guaranteed to control the Type I error under the null hypothesis when the two distributions are equal (see e.g., Lehmann and Romano, 2005b).
@@
Results: Evaluating power. Figure˜16 presents the performance of different hypothesis testing methods as a function of the real sample size $n$, for $N=100$, $\alpha=10\%$, and $\varepsilon=5\%$. The rejection rate (power) is well above the target level $\alpha$, indicating that the mean activation score for images containing animals is indeed higher than for non-animals, supporting the conclusion that attention head L22H6 detects animals.
Figure 16: Performance comparisons as a function of the real sample size $n$. Hypothesis testing methods applied to animal versus non-animal groups, each containing three classes, at target level $\alpha=10\%$ and $\varepsilon=5\%$.
Following that figure, we observe that GESPI consistently achieves higher power than the base method OnlyReal across all values of $n$, with power increasing as $n$ grows, illustrating the benefit of leveraging additional synthetic data. For sufficiently large $n$ ($n\geq 45$), both methods achieve power close to $1$ and obtain comparable performance. At the same time, OnlySynth, which relies only on synthetic images, achieves high power due to the large sample size; however, this approach does not provide valid error rate control, as the synthetic data may differ from the real data distribution.
Results: Evaluating Type I error. Next, we compare the performance of different hypothesis testing methods under the null hypothesis. To do so, we split the non-animal classes into two disjoint groups and repeat the experiment from Figure˜16. Under the null, the activation score distributions of the two groups are expected to be equal. Indeed, Figure˜17(a) shows the histograms for the real data, where the two groups exhibit similar distributions. In contrast, Figure˜17(b) shows the histograms for the synthetic data, where the two groups display noticeable differences, highlighting that the synthetic distribution does not perfectly match the real one.
@@
Figure 17: Histograms of the per-image activation scores $s^{l,h}(X)$ for two disjoin non-animal groups. Each histogram is based on 1,000 samples drawn from the corresponding group; (a) shows results for real ImageNet images, and (b) for synthetic images generated with FLUX.1.
Figure˜18 shows the Type I error for $n=10$ and $100$; other details are as in Figure˜16. Both OnlyReal and GESPI control the Type I error at the target level $\alpha$. In contrast, OnlySynth obtains very high Type I error, illustrating that the synthetic data cannot be naively treated as if it were real.
@@
Figure 18: Performance comparisons under the null hypothesis. Hypothesis testing methods applied to two disjoint groups of non-animal classes at target level $\alpha=10\%$ and $\varepsilon=5\%$; with (a) $n=10$ and (b) $n=100$.
@@ Appendix D Examples of statistical inference with risk control
Table 2: Summary of examples of our framework.
@@ Appendix F Additional theoretical results for GESPI
The first two statements in Condition F.1 state that there exists a partial order on the action space with “reasonable properties”—namely, that the minimum and maximum of any two elements are well-defined and that the distributive law holds.
Table 3 illustrates what the partial order means in our examples. The monotonicity of the algorithm and loss can be checked on a case-by-case basis. For instance, conformal prediction sets are nested by construction (Vovk et al., 2005). The requirements in Condition F.1 hold in a variety of statistical inference problems, including the examples in Section 3. See also Table 3 for simple examples.
@@
Table 3: Examples of partially ordered action spaces for different inference problems. Here, $\mathcal{B}(\mathcal{X})$ denotes the Borel sigma-algebra of $\mathcal{X}$.

### v2

@@ C.3 Conformal Risk Control for Protein Structure Prediction
Figure˜S11 visualizes protein T1078, for which AlphaFold produces a highly accurate predicted structure. Figures˜11(a) and 11(b) present the resulting prediction sets obtained by OnlyReal and GESPI, respectively. Panel (1) presents the predicted structure, with residues abstained on marked in red and accepted residues in blue. Both methods obtain risk equal to 0; however, OnlyReal conservatively abstains from most residues ($\sim 91\%$), while GESPI abstains from significantly fewer ($\sim 48\%$), demonstrating the benefit of leveraging synthetic data. Panel (2) shows the predicted structure (red and blue) aligned with the true structure (gray), illustrating that most of the protein is accurately predicted. Panel (3) highlights a small subset of residues whose predicted structure is inaccurate; both methods abstain. Lastly, panel (4) shows a well-predicted region where OnlyReal unnecessarily abstains, while GESPI correctly accepts the residues.
@@
Figure S11: Visualization of protein structure prediction with error rate control. Panels show protein T1078 predictions with residues abstained on by (a) OnlyReal and (b) GESPI methods. Red: residues abstained on; Blue: accepted residues. Gray: real experimental structure, aligned with AlphaFold2 predicted structure. Quantitative results {abstention ratio, risk}: OnlyReal – $\approx 91\%$, $0\%$}; GESPI – $\approx 48\%$, $0\%$}.
Figure˜S12 complements Figure˜4 in the main manuscript by showing the chosen pLDDT thresholds, $\hat{\lambda}$, for $\alpha=5\%$ and $15\%$. The figure demonstrates that the selected thresholds vary with $\alpha$ and differ across methods: OnlyReal conservatively chooses higher thresholds, while our GESPI procedure selects lower thresholds with lower variability than Guardrail. As a result, GESPI achieves risk closer to the nominal $\alpha$ level with lower variability.
@@
Figure S12: Chosen pLDDT thresholds, $\hat{\lambda}$, for $\alpha=5\%$ and $15\%$, complementing the results shown in Figure˜4.
@@
Figure˜S13 presents the performance of the conformal risk control methods at target level $\alpha=10\%$. The results show a similar trend to that in Figure˜4. In particular, OnlyReal conservatively controls the risk and leads to a higher and more variable abstention rate. OnlySynth, which relies on approximate prediction errors, does not provide valid risk control guarantees. In contrast, our proposed method, GESPI, achieves risk close to the nominal $\alpha$ level with a lower abstention rate and lower variability compared to Guardrail.
@@
Figure S13: Performance comparisons for protein structure prediction with error rate control. Conformal risk control methods applied at target level $\alpha=10\%$. Left: average risk (fraction of residues with error >3>3Å). Middle: average abstention rate. Right: selected pLDDT threshold $\hat{\lambda}$.
C.4 Hypothesis Testing for Comparing Large Reasoning Models
For brevity, we provide an overview of the experimental setup here and refer to Section˜B.3 for full details.
Given two large language models (LLMs), we aim to test whether model A outperforms model B on a specific type of task. Formally, we consider the hypotheses $\mathcal{H}_{0}:p\leq 0.5\text{ vs. }\mathcal{H}_{1}:p>0.5,$ where $p$ denotes the win rate of model A over model B.
Data. We evaluate the win rate on the AIME25 dataset, which consists of 30 challenging reasoning math questions. The goal is to pin down the ranking of the models on AIME25 as closely as possible, which is hindered by the very small number of questions for this competition. For synthetic data, we use a subset of the OlympiadBench (He et al., 2024) math competition questions, which resemble AIME problems but are drawn from a different distribution and therefore cannot be treated as real test data. [^9] Since OlympiadBench was released in 2024, before AIME25, these two datasets are non-overlapping. For each question and model, we record whether the model’s answer is correct. Further details on the experimental setup are provided in Section˜B.3.
We consider two experiments: the first is used for hyperparameter selection in the K-way majority vote, and the second compares the performance of different models.
C.4.1 Hyperparameter Selection for K-Way Majority Vote
Majority-vote scheme typically improves performance as $K$ increase, but at the cost of requiring $K$ forward passes per input. In this experiment, we compare the performance of the R1-Deepseek-distill-Qwen-7B model with an expensive majority vote using $K=64$ to cheaper alternatives with $K=1,4,8,$ and $16$.
Figure˜S14 presents the performance of different hypothesis testing methods across these different $K$-way majority votes. All methods control the type I error close to the nominal level $\alpha$, except Guardrail, which always controls at the higher level $\alpha+\varepsilon$ by design. Note that OnlySynth performs the test on a different dataset, which is intended to be similar to the real datset (AIME25) but may not follow the same distribution; therefore, it is not a competing method and is included for illustration only.
Following the left panel, we see that comparing $K=64$ to $K\in\{1,4,8\}$ yields rejection rates above the nominal level, indicating that the alternative holds, with GESPI achieving higher power than OnlyReal and is comparable with Guardrail. Interestingly, there is no strong evidence that $K=64$ outperforms $K=16$ for this task, as all methods achieve rejection rates close to $\alpha$.
Figure S14: Performance comparisons for different K-way majority vote on the AIME25 dataset for the R1-Deepseek-distill-Qwen-7B model. Hypothesis testing methods are applied at level $\alpha=5\%$ and $\varepsilon=2\%$. Comparisons are made between the expensive $K=64$ and cheaper $K=1,4,8,16$. Left: Rejection rate, measured under the standard setting. Right: Type I error, measured in the shuffled-response setting where the null hypothesis holds.
C.5 Comparing Different Models
Figure˜S15 presents two model comparisons, each involving a distinct pair of LLMs. In both cases, the rejection rate (power) is well above the target level $\alpha$, indicating that the first model outperforms the second. Our proposed method, GESPI, achieves higher power than OnlyReal and is comparable with Guardrail. In the shuffled-answers setting, both GESPI and OnlyReal methods achieve Type I error at the target level $\alpha$, while Guardrail controls the error at the higher level $\alpha+\varepsilon$.
Figure S15: Performance comparisons for LLM win rate on AIME25 dataset. Hypothesis testing methods applied at level $\alpha=5\%$ and $\varepsilon=2\%$. Left: Description of model comparisons. Middle: Power, comparing the rejection rate under the standard setting. Right: Type I error, measured in the shuffled-response setting where the null holds.
C.6 Single and Multiple Hypothesis Testing for Outlier Detection
We provide results for single and multiple hypothesis testing for outlier detection at specific $\alpha$ and $\varepsilon$ levels, representing a real-world scenario, as well as ROC curves for different trimming proportions. We note that the ROC curves shown in the main manuscript require knowledge of the true outliers and inliers—information that is not available in practice—and are therefore included only for illustrative comparison.
Figures˜S16 and S17 present the performance for single and multiple testing, for both $q=2.5\%$ and $5\%$, respectively. In both cases, we observe similar trends. OnlyReal and Guardrail suffer from the small sample size: the former conservatively controls the error rate but obtains low power, whereas the latter controls the error rate at a higher level while still exhibiting high variability. The OnlySynth produces arbitrary error rates, while the Oracle method achieves error rate tightly regulated around the target level, as expected. In contrast, the error rate of GESPI is close to the nominal $\alpha$ level, while achieving substantially higher power than OnlyReal, lower variability than Guardrail, and performance approaching that of the Oracle.
@@ C.6 Single and Multiple Hypothesis Testing for Outlier Detection
Figure S16: Performance comparisons for outlier detection. Evaluated on three datasets: (1) shuttle, (2) credit-card, (3) KDDCup99. Left two panels: single-outlier case ($\alpha=2\%$, $\varepsilon=1\%$). Right two panels: multiple-outlier case ($\alpha=15\%$, $\varepsilon=10\%$). The trimming proportion is $q=2.5\%$.
Figure S17: Performance comparisons for outlier detection. Evaluated on three datasets: (1) shuttle, (2) credit-card, (3) KDDCup99. Left two panels: single-outlier case ($\alpha=2\%$, $\varepsilon=1\%$). Right two panels: multiple-outlier case ($\alpha=15\%$, $\varepsilon=10\%$). Same setup as in Figure˜S16, but with $q=5\%$ trimming proportion.
Figure˜S18 presents the empirical power of multiple testing methods as a function of the empirical FWER for trimming proportion $q=5\%$, complementing the main manuscript results for $q=2.5\%$. For a trimming proportion of $5\%$, where the synthetic data are of lower quality, the ROC curve of GESPI closely follows that of OnlyReal. Notably, even in this regime, GESPI still exhibits lower variability in the error rate, as observed in Figure˜S17.
Figure S18: Empirical power versus empirical FWER for outlier detection. Results are shown for trimming proportion $5\%$, corresponding to a scenario with lower quality synthetic data. All other details follow in Figure˜S16 for the FWER experiment.
C.7 Hypothesis Testing for Mechanistic Interpretability of a Vision Transformer Model
@@
In this section, we consider a task related to the mechanistic interpretability of vision transformers. We aim to test whether a specific attention head exhibits a functional role—for example, detecting shapes, animals, or spatial locations. Concretely, we test whether the activation of a given head is stronger when the object it is hypothesized to detect is present in the image compared to when it is absent, which would provide evidence that the head indeed serves this role. As a case study, we examine the attention head at layer 22, head 6 (L22H6) in CLIP ViT-L/14 (Radford et al., 2021b), which has been previously reported to strongly respond to images of animals (Gandelsman et al., 2024).
@@
We use ImageNet images as the real dataset, partitioned into animals and non-animals classes. For the synthetic dataset, we generate images using FLUX.1 (Black Forest Labs, 2024), also separated into animals and non-animals. Details for each dataset are as follows:
@@
  • FLUX.1 (synthetic) (Black Forest Labs, 2024): For each class, we generate 2,0002{,}000 images using the FluxPipeline from diffusers with mixed precision (float16) and $50$ inference steps. Prompts follow the CLIP-style template “A photo of a {class name}” (Radford et al., 2021b), where {class name} is the corresponding ImageNet label. The resulting synthetic images are labeled animals / non-animals based on the originating class.
@@
Intuitively, $H^{l,h}_{i}(X)$ measures how strongly token $i$ aligns with the animal text direction. See an example visualized as a spatial heatmap in Figure S19.
Figure S19: Left: input image $X$ of a dog. Right: spatial heatmap $H^{22,6}(X)$ from layer 22, head 6, obtained by projecting token contributions onto $M_{\text{text}}(\text{``a photo of an animal''})$.
@@
Formulation of the null hypothesis. We compute the activation score $s^{l,h}(X)$ (7) on two disjoint image sets, $\mathcal{D}_{\text{animals}}$ and $\mathcal{D}_{\text{non-animals}}$, and test whether head $(l,h)$ exhibits the same activation on animal and non-animal images. Let $P^{(l,h)}_{\text{animals}}$ and $P^{(l,h)}_{\text{non-animals}}$ denote the population distributions of activation scores of head $(l,h)$ over animals and non-animals, respectively. We test the null hypothesis that these two distributions are equal:
$\mathcal{H}_{0}:P^{(l,h)}_{\text{animals}}=P^{(l,h)}_{\text{non-animals}}$.
@@
To set the critical value, we use a standard permutation approach, which is guaranteed to control the Type I error under the null hypothesis when the two distributions are equal (see e.g., Lehmann and Romano, 2005b). We use a one-sided test, assigning smaller $p$-values to larger values of $\hat{\Delta}^{(l,h)}$, corresponding to stronger activation on animal images relative to non-animal images.
@@
Results: Evaluating power. Figure˜S20 presents the performance of different hypothesis testing methods as a function of the real sample size $n$, for $N=100$, $\alpha=10\%$, and $\varepsilon=5\%$. The rejection rate (power) is well above the target level $\alpha$, indicating that the mean activation score for images containing animals is indeed higher than for non-animals, supporting the conclusion that attention head L22H6 detects animals.
Figure S20: Performance comparisons as a function of the real sample size $n$. Hypothesis testing methods applied to animal versus non-animal groups, each containing three classes, at target level $\alpha=10\%$ and $\varepsilon=5\%$.
Following that figure, we observe that GESPI consistently achieves higher power than the base method OnlyReal across all values of $n$, with power increasing as $n$ grows, illustrating the benefit of leveraging additional synthetic data. For sufficiently large $n$ ($n\geq 45$), both methods achieve power close to $1$ and obtain comparable performance. For all values of $n$, GESPI and Guardrail attain the same power. At the same time, OnlySynth, which relies only on synthetic images, achieves high power due to the large sample size; however, this approach does not provide valid error rate control, as the synthetic data may differ from the real data distribution.
Results: Evaluating Type I error. Next, we compare the performance of different hypothesis testing methods under the null hypothesis. To do so, we split the non-animal classes into two disjoint groups and repeat the experiment from Figure˜S20. Under the null, the activation score distributions of the two groups are expected to be equal. Indeed, Figure˜21(a) shows the histograms for the real data, where the two groups exhibit similar distributions. In contrast, Figure˜21(b) shows the histograms for the synthetic data, where the two groups display noticeable differences, highlighting that the synthetic distribution does not perfectly match the real one.
@@
Figure S21: Histograms of the per-image activation scores $s^{l,h}(X)$ for two disjoin non-animal groups. Each histogram is based on 1,000 samples drawn from the corresponding group; (a) shows results for real ImageNet images, and (b) for synthetic images generated with FLUX.1.
Figure˜S22 shows the Type I error for $n=10$ and $100$; other details are as in Figure˜S20. OnlyReal, Guardrail, and GESPI all attain Type I error below the target level $\alpha$. In contrast, OnlySynth obtains very high Type I error, illustrating that the synthetic data cannot be naively treated as if it were real. Since the synthetic data do not follow the null in this setting, as further illustrated in Figure˜S21, GESPI attains a Type I error rate similar to Guardrail.
@@
Figure S22: Performance comparisons under the null hypothesis. Hypothesis testing methods applied to two disjoint groups of non-animal classes at target level $\alpha=10\%$ and $\varepsilon=5\%$; with (a) $n=10$ and (b) $n=100$.
@@ Appendix D Examples of Statistical Inference with Risk Control
Table S1: Summary of examples of our framework.
@@ Appendix F Additional Theoretical Results for GESPI
The first two statements in Condition F.1 states that there exists a partial order on the action space with “reasonable properties”—namely, that the minimum and maximum of any two elements are well-defined and that the distributive law holds.
Table S2 illustrates what the partial order means in our examples. The monotonicity of the algorithm and loss can be checked on a case-by-case basis. For instance, conformal prediction sets are nested by construction (Vovk et al., 2005). The requirements in Condition F.1 hold in a variety of statistical inference problems, including the examples in Section 3. See also Table S2 for simple examples.
@@
Table S2: Examples of partially ordered action spaces for different inference problems. Here, $\mathcal{B}(\mathcal{X})$ denotes the Borel sigma-algebra of $\mathcal{X}$.